OpenStudy (anonymous):

How long will it take an investment to double in value if the interest rate is 8% compounded continuously? What is the equivalent annual interest rate?

OpenStudy (anonymous):

want the investment to double, 8% continously A=Pe^rt, where A=final amount,P=principal,r=rate,t=years 2P=Pe^rt, 2=e^rt, ln(2)=rt t=ln(2)/(8/100)=---?

OpenStudy (anonymous):

@ Aenn88,let me know if you have question :D